A toy truck travels a distance of 2.5 meters in 0.95 seconds after a ball is shot out from it in the opposite direction. The mass of the truck is 1kg angmdthe ball is 0.06 kg. What is the velocity of the ball.


1
Expert's answer
2021-01-19T07:09:42-0500

First, find the initial speed of the toy truck:


"v=2d\/t."


Next, according to conservation of momentum, we can write


"Mv=mu, \\\space\\
u=\frac{Mv}{m}=\frac{2dM}{mt}=\frac{2\cdot2.5\cdot1}{0.06\cdot0.95}=87.7\text{ m/s}."